Shift report 20000822 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Aug 22 2000 - 00:01:48 EDT

  • Next message: http account: "Shift report 20000822 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Aug 22 00:01:47 2000, Shift: 16:00-24:00
    Supervisor: FR, with: Ronny
    Log book#4 pages 137 - 137
    
    Shift Summary:
    ------------------------------------------------------------
     1. No beam during entire shift. They started to inject
      into RHIC at 22:00 but all attempts failed so far.
    
     2. No data taken.
      T1 anode HV is ok, leakage current around 15nA.
    
     3. D1, D2 and D3 have been off all shift.
        All TPC anode voltages off. 
        No problems encountered during shift check.
       
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Aug 22 2000 - 00:05:21 EDT